Hi everyone
My problem is:
Give exact and approximate solutions to three decimal places
x^2+6x+9=25
Thanks you
Gina

Answer by ankor@dixie-net.com(22740)

My problem is:
Give exact and approximate solutions to three decimal places
x^2 + 6x + 9 = 25
x^2 + 6x + 9 - 25 = 0
x^2 + 6x - 16 = 0
This factors exactly to
(x+8)(x-2) = 0
Two solutions
x = -8
and
x = +2
:
An approximate solution would not be necessary.